OpenStudy (anonymous):

what should be added to both sides of this equation to solve for the variabel. t+(-8)=4

OpenStudy (anonymous):

8

OpenStudy (anonymous):

t-8+8=4+8 t+(-8)=4 t=12 t-8=4 t=4+8 t=12

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

but should i type th answer as 12 or 8?

OpenStudy (anonymous):

@cornitodisc

OpenStudy (anonymous):

the question is what number should be added,,so its 8

OpenStudy (anonymous):

thanks